# Investigating Large Language Models in Diagnosing Students’ Cognitive Skills in Math Problem-solving

### v1

@@ 5 Results
In this section, we organize and interpret the findings to answer our research questions. The complete results from our experiment are presented in Table 6 in the Appendix.
5.1 RQ1. How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
Overall, the results underscore the difficulty of the task (Table 2); no model surpassed an F1 score of 0.5, indicating that even state-of-the-art models struggle to diagnose students’ cognitive skills accurately. Our manual error analysis on a few samples revealed that LLMs largely failed when 1) student response flows in a non-directional, scattered layout, 2) a student uses underscores, strikethroughs, and calculation notes, and 3) a response lacks direct evidence for a skill, despite contextual information being sufficient for inference.
[…]
Models Precision Recall F1 Accuracy Overconfidence Underconfidence
DeepSeek-V3 .406 .461 .417 .714 .818 .107
GPT-4o .398 .476 .412 .672 .788 .110
Llama-3.1-405B .415 .438 .384 .624 .778 .186
Claude-3.5-Sonnet .329 .387 .332 .654 .812 .173
Gemini-1.5-Pro .430 .496 .441 .707 .840 .132
Table 2: The performance of the five state-of-the-art LLMs on the cognitive skill diagnosis task.
Analysis of skill-specific performance (Figure 2) reveals further insights into the limitations of current LLMs. No skill category achieved a maximum F1 score above 0.5. Interestingly, the skill Compute, which most closely resembles conventional grading tasks (i.e., evaluating the correctness of mathematical procedures and calculations), still yielded only a moderate F1 score across models. This moderate F1 score confirms LLMs’ struggle in verifying correct computation in constructed student responses.
Contrary to expectations, there was no substantial performance gap between the “Knowing” and “Applying” cognitive skills. Although we anticipated that “Knowing” skills (e.g., Recall, Recognize) would be easier than “Applying” skills for LLMs due to their surface-level nature, models performed comparably. An exception is found in the skill Classify/Order, which shows a relatively high average accuracy but a lower F1 score. Closer inspection reveals that “Evident Yes” is dominant in the ground truth dataset (Table 5), leading to the discrepancy between F1 score and accuracy.
Figure 2: The performance of five state-of-the-art LLMs across cognitive skill types. The left plot shows the minimum (bottom $\times$), maximum (top $\times$), and average (middle •) F1 scores for each skill category, while the right plot presents the corresponding accuracy metrics.
5.2 RQ2. How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
We observed subtle impacts of multimodal input and reasoning on the performance, while the impact of model size was significant.
Multimodal Input. Models supporting image input (e.g., GPT-4o-img, Claude-3.5-Sonnet-img, Gemini-1.5-Flash-img) consistently outperformed their text-only counterparts in accuracy (up to 10.6%) and in reducing underconfidence (up to 4.9%) (Figure 3 left), although the statistical difference between the pairs was not significant. Image input was particularly beneficial in cases where OCR transcription introduced errors. For example, a student underlined an arithmetic expression and wrote the corresponding calculation beneath it (see Figure 5 a). The OCR system misinterpreted this layout as a fraction, leading Claude-3.5-Sonnet (text-only) to incorrectly conclude that the student had set up an invalid equation. In contrast, the same model with image input correctly inferred the student’s intention and produced an accurate diagnosis.
Reasoning. The performance of reasoning-oriented models showed mixed results (Figure 3 right). DeepSeek-R1 achieved the highest F1 and accuracy scores among all models. However, other reasoning models, such as o1-Preview and Gemini-2.0-Flash-Thinking, did not show clear advantages over standard models in either accuracy or F1 score. These results suggest that reasoning capability alone does not guarantee better performance. For instance (Figure 5 b), DeepSeek-R1 classified a response as “Vague No” because the student omitted the factorization steps. However, the original expression was $\sqrt{48a}$, and the model failed to detect a clear error in the variable term, indicating that the student’s reasoning was evidently flawed. Throughout its reasoning process, the model focused on the constant term only and overlooked other possible places of error. While current reasoning models excel at pushing forward linear, step-by-step logic, effective cognitive diagnosis may also require divergent, multi-perspective reasoning for thorough analysis.
[…]
Figure 3: Impact of multimodal input (left) and reasoning capability (right) on LLMs’ performance in cognitive skill diagnosis. On the left, each adjacent pair of bars shows performance with and without the image input from the same model family. On the right, models with reasoning capabilities are grouped at the front.
Figure 4: Impact of model size on LLM performance in cognitive skill diagnosis. Models are grouped into large, medium, and small categories.
@@ 6 Discussion
Our findings imply that current LLMs are not yet suitable for high-stakes cognitive skill diagnosis. The models exhibit low recall and high overconfidence (Table 2), raising concerns that LLM-based systems may frequently overlook students’ misconceptions. This shortcoming is particularly problematic given that the primary goal of automated assessment is to identify and address such misconceptions, and it can compromise teachers’ and students’ trust in automated systems (Shin, 2021). To responsibly integrate LLMs into real-world assessment settings, practitioners should begin by evaluating their performance on small, ground-truth samples. This process not only exposes limitations but also allows for fine-tuning model judgments and confidence levels for downstream math topics and student populations. In parallel, model developers may work to mitigate overconfidence during training by encouraging more conservative, evidence-based decision-making.
Our manual analysis of failed cases reveals that inferring students’ thought processes from handwritten responses presents distinct challenges. While mathematical problem-solving is inherently procedural, students often express their reasoning in non-linear, spatially scattered layouts that lack clear directionality. This ill-structured presentation differs significantly from the format of most existing reasoning and problem-solving benchmarks. Moreover, ambiguous notations (e.g., fractions misinterpreted as underscores) and unconventional layouts frequently result in OCR transcription errors, which hinder accurate diagnosis unless image inputs are provided. To enable more reliable inference, future work may require richer representations of math responses that capture high-level semantic intent beyond surface-level positional and textual cues.
7 Limitations and Future Work
While MathCogrepresents a novel benchmark for evaluating LLMs on cognitive skill diagnosis, it has several limitations. First, due to data-sharing restrictions from the source provider (AIHub), the original math problems and student responses cannot be publicly released outside Korea. To support future research despite this constraint, we release the complete set of diagnostic checklists and expert-generated human annotations, which can serve as references for researchers to develop similar datasets in their educational contexts. In parallel, we are actively collecting new student response data under informed consent, with the goal of constructing a fully public benchmark in the future.
This ongoing data collection effort will also address limitations in the diversity of problem types represented in MathCog. Currently, the dataset focuses on nine problem types, primarily centered around arithmetic equation solving. Other areas of mathematics—such as geometry—often require interpreting visual representations and may further benefit from the integration of multimodal input. Additionally, reasoning-oriented problems, which explicitly ask students to generalize mathematical knowledge or justify their arguments, are underrepresented. We plan to expand the dataset to include a broader range of content areas and grade levels, enabling a more holistic evaluation of LLM performance across the mathematics curriculum.
Our experimental results are also bound to zero-shot settings, without exploring the effects of few-shot or test-time compute (Snell et al., 2024) prompting. Future work can investigate whether more sophisticated prompting enhances diagnostic accuracy. For example, prompting an LLM to explicitly infer a student’s thinking process before answering each check item may lead to more accurate judgments. We see our work as a first step toward automated cognitive skill assessment, and we invite the community to build upon these directions.
Acknowledgments
This work was supported by Algorithm LABS. This research (paper) used datasets from “The Open AI Dataset Project (AI-Hub, S. Korea).” All data information can be accessed through “AI-Hub (www.aihub.or.kr).”
Appendix A Appendix
@@
Carry out algorithmic procedures for +,,×,÷+,-$\times$,$\div$+ , - , × , ÷ or a combination of these
@@
Problem Types 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
@@
Table 4: The inter-rater percentage absolute agreement of each problem type. The percentage indicates the ratio of unanimous verdicts in each teacher group.
A.1 Prompts
A.2 System Prompt
# **Task Description**
## **Evaluation Categories**
## **Input Format**
You will receive the following data:
- **Problem**: A math problem given to a student.
- **Answer**: The correct step-by-step solution.
- **Response**: The students response to the problem.
- **Check Items**: A set of specific skills to evaluate.
## **Output Format**
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
A.3 User Prompt
# **Task**
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
**Problem:**
There is a two-digit natural number whose tens digit is $1$. If the number that changes the tens and ones digits of this natural number is $9$ less than $5$ times the first number, find the first number.
**Answer:**
If the number in the ones place is $x$, this natural number is $10+x$. The number where the tens digit and the ones digit are swapped is $110x+$ 1 because the tens digit is $x$ and the ones digit is $1$. The changed number is $9$ smaller than $5$ times the first number, so $910x+1=5(10+x)-$ 9 $4110x+1=5x+$ 41 $405x=$ 40 $x=8$ Therefore, the first number is $18$.
**Response:**
$1$ $x$ $x$ $1$ $910x+1=5(10+x)-$ 9 $910x+1=50+5x-$ 9 $41$ $405x=$ 40 $x=8$
**Check Items:**
**Output:**
‘‘‘json
"skills": [
"evidence": "10 x+1=5(10+x)-9",
"verdict": "Evident Yes"
"evidence": "10 x+1=5(10+x)-9",
"verdict": "Evident Yes"
Recall Recognize Classify/Order Compute Retrieve Determine Represent Implement
Evident Yes 507 506 92 584 9 441 154 319
Evident No 78 69 8 160 9 54 55 146
Vague Yes 58 54 2 48 0 25 6 29
Vague No 98 62 0 58 34 16 46 41
Student responses 582 583 51 796 52 482 261 535
@@ A.3 User Prompt
Models Precision Recall F1 Accuracy Overconfidence Underconfidence
DeepSeek-R1 .447 .467 .442 .773 .901 .028
DeepSeek-V3 .406 .461 .417 .714 .710 .107
GPT-4o-img .427 .494 .448 .743 .738 .070
GPT-4o .398 .476 .412 .672 .703 .110
GPT-4o-Mini .345 .397 .352 .622 .791 .102
o1-Preview .338 .410 .354 .711 .757 .117
Llama-3.3-70B .416 .468 .400 .640 .534 .189
Llama-3.1-8B .283 .265 .257 .642 .695 .138
Llama-3.1-405B .415 .438 .384 .624 .451 .186
Claude-3.5-Sonnet-img .326 .365 .330 .690 .680 .140
Claude-3.5-Sonnet .329 .387 .332 .654 .676 .173
Gemini-1.5-Flash-img .431 .484 .429 .702 .593 .129
Gemini-1.5-Flash-8b .287 .341 .279 .558 .579 .196
Gemini-1.5-Flash .432 .490 .432 .679 .582 .155
Gemini-1.5-Pro .430 .496 .441 .707 .721 .132
Gemini-2.0-Flash-Thinking .455 .492 .421 .690 .589 .186
Mistral-7B .384 .369 .347 .714 .805 .112
Qwen-2.5-7B .278 .270 .271 .705 .842 .075
Qwen-2.5-72B .415 .459 .416 .710 .631 .115
Table 6: Performance comparison of all 16 LLMs tested on skill diagnosis tasks.

### v2

@@ 5 Results
This section reports results addressing our research questions.
Category Model Precision Recall F1 Accuracy OverAttr FalseAttr
Reasoning Model DeepSeek-R1 .447 .467 .442 .773 .767 .862
Gemini-2.0-Flash-Thinking .454 .486 .418 .688 .361 .432
o1-Preview .423 .513 .443 .711 .489 .650
Multimodal Model Claude-3.5-Sonnet-img .408 .457 .413 .691 .626 .559
Gemini-1.5-Flash-img .431 .484 .429 .702 .471 .471
GPT-4o-img .427 .494 .448 .743 .656 .681
Large Model Claude-3.5-Sonnet .411 .487 .416 .656 .551 .574
DeepSeek-V3 .406 .461 .417 .714 .612 .595
Gemini-1.5-Pro .429 .496 .440 .706 .520 .635
GPT-4o .398 .476 .412 .672 .648 .658
Llama-3.1-405B .415 .439 .385 .624 .467 .307
Medium Model Llama-3.3-70B .415 .467 .397 .635 .436 .409
Qwen-2.5-72B .415 .459 .416 .711 .568 .500
Small Model Gemini-1.5-Flash .432 .490 .432 .679 .502 .465
Gemini-1.5-Flash-8b .359 .427 .348 .558 .537 .525
GPT-4o-mini .345 .398 .352 .622 .767 .784
Llama-3.1-8B-128K .352 .334 .323 .653 .709 .624
Qwen-2.5-7B .347 .337 .339 .705 .758 .798
Table 2: Performance of all 18 LLMs tested on skill diagnosis tasks. For each metric, blue and red fonts indicate the best and worst values, respectively.
Figure 2: F1 score across by model type and family.
5.1 RQ1. How do different LLMs perform in diagnosing students’ cognitive skills?
Overall, all evaluated LLMs showed limited performance in cognitive skill diagnosis, with macro F1 scores below 0.5 (Figure 2, Table 2). While overall accuracy was relatively moderate ($M=.680$, $SD=.050$), it obscures important diagnostic failures. Models frequently over-attributed evidential strength (OverAttr; $M=.580$, $SD=.120$) and falsely attributed evidence (FalseAttr; $M=.585$, $SD=.145$), often asserting strong evidence even when diagnoses were incorrect. Notably, even the best-performing models in terms of F1 score and accuracy (e.g., DeepSeek-R1, GPT-4o-img) showed high OverAttr and FalseAttr values, suggesting that performance gains often co-occur with unreliable evidential judgement.
Figure 3: Performance and evidential attribution errors across cognitive skills, measured by F1 score, accuracy, evidence over-attribution, and false-attribution.
Analysis of skill-specific performance (Figure 3) reveals further insights into the limitations of current LLMs. No skill category achieved an F1 score above 0.5, indicating that fine-grained diagnosis remains challenging even at the individual skill level. Despite this overall limitation, we observed a performance gap between the Knowing and Applying cognitive skills. Contrary to our expectation that Knowing skills (e.g., Recall, Recognize) would be easier due to their surface-level nature, models performed better on Applying skills in both F1 score ($t=4.34$, $p<.001$) and accuracy ($t=4.85$, $p<.001$). However, this performance gain came with increased evidential errors. Applying skills exhibited higher evidence over-attribution ($t=5.44$, $p<.001$) and false-attribution ($t=7.66$, $p<.001$) than Knowing skills, indicating a tendency to overstate evidential support when diagnosing higher-level cognitive skills.
Few-shots Precision Recall F1 Accuracy OverAttr FalseAttr
claude-3-5-sonnet .398 .485 .410 .646 .515 .647
deepseek-V3 .404 .493 .404 .624 .493 .497
gemini-1.5-pro .429 .517 .430 .665 .427 .552
gpt-4o .390 .485 .396 .637 .577 .587
llama-3.1-405B .442 .494 .429 .677 .476 .457
Table 3: Performance of five state-of-the-art models in few-shot experiments. Bold values denote improvements over the corresponding zero-shot results.
We further examined the impact of multimodality, reasoning, model size, and few-shot prompting as factors that may influence diagnostic performance (Table  2). Multimodal input provided modest but inconsistent benefits. Models supporting image input consistently outperformed their text-only counterparts in accuracy, indicating that access to handwritten layouts and visual cues helps mitigate errors introduced by OCR transcription. However, improvements in F1 score were inconsistent, suggesting that visual input can sometimes introduce noise and does not uniformly improve fine-grained skill diagnosis. Reasoning-oriented models produced similarly mixed results. While some reasoning-oriented models (e.g., DeepSeek-R1, GPT-o1) achieved higher F1 scores and accuracy than their non-reasoning counterparts, Gemini performed worse. These results suggest that reasoning mechanisms do not uniformly translate into improved cognitive diagnosis. Model size showed a moderate positive relationship with performance. Larger models generally achieved higher F1 scores than smaller ones, indicating that increased parameter capacity supports a more robust interpretation of student responses. However, the correlation was moderate (Spearman’s $\rho=.570,p=.053$), suggesting that scale alone is insufficient to account for the substantial variability in performance across models. Finally, few-shot prompting did not improve diagnostic performance. While in-context examples might seem helpful for such nuanced tasks, F1 and accuracy often decreased rather than improved (Table 3), suggesting the challenge lies in the complexity of cognitive skill diagnosis rather than prompt design. However, few-shot prompting did reduce tendencies toward evidence over-attribution and false-attribution, indicating that examples may encourage more cautious, evidence-grounded responses even when overall diagnostic accuracy remains limited. Taken together, these findings suggest that prompt engineering and model capabilities have limited impact on improving the performance of cognitive skill diagnosis.
5.2 RQ2. How does the evidential strength of student responses affect LLM’s cognitive skill diagnosis performance?
Figure 4: Effect of evidential strength on LLM diagnostic performance for every model.
Our results reveal a strong dependency of LLM diagnostic performance on the evidential strength of student responses. Across all models, both F1 score ($t=64.60$, $p<.001$) and accuracy ($t=17.80$, $p<.001$) were substantially higher when student responses contained evident evidence for a cognitive skill than when evidence was vague (Fig. 4). The magnitude of this gap was substantial (mean $\Delta F1=.51$; mean $\Delta Acc=.41$), indicating that current LLMs rely heavily on explicit evidence and exhibit limited robustness when reasoning under weak or implicit evidential conditions.
Beyond degraded performance, models also exhibited over-attribution bias under vague evidence. Specifically, they frequently labeled responses as Evident even when the ground-truth evidential strength was Vague, resulting in consistently high OverAttr values. More concerningly, models also showed a strong tendency toward FalseAttr, in which they asserted evident evidence despite arriving at incorrect cognitive skill diagnoses. In such cases, models produced unsupported explanations that appeared plausible yet were not grounded in students’ actual responses. Notably, these two tendencies were strongly correlated across models (Spearman’s $\rho=.837$, $p<.001$), indicating that models prone to over-attributing evidential strength are also more likely to produce incorrect diagnoses with asserted evidence. This behavior is particularly problematic for educational use, as it can mislead teachers and students by presenting incorrect diagnoses with seemingly well-justified rationales [15].
Error code Description Count (%)
E1. Evidence Misidentification Incorrectly identifying or relying on insufficient evidence from the student response. 659 (23.40%)
E2. Rubric Misinterpretation Misunderstanding the diagnostic rubric or evaluation criteria. 279 (9.91%)
E3. Over-Inference Inferring unstated reasoning beyond the information provided in the student response. 942 (33.45%)
E4. Consistency Issue Inconsistencies between evidence, explanation, and final verdict. 379 (13.46%)
E5. Hallucination Introducing content or reasoning not present in the student response. 557 (19.78%)
Table 4: Error types and their distribution identified from qualitative analysis of LLM diagnostic failures under vague evidence.
To better understand how such errors arise, we conducted a qualitative error analysis of model outputs under incorrectly diagnosed vague evidence cases. We examined all components generated in the models’ CoT responses, including the identified evidence, explanations, and final verdicts, focusing on how the models misinterpreted students’ answers, generated unsupported reasoning, and ultimately reached incorrect diagnoses. Three authors collaboratively analyzed 372 cases (13.2%) sampled from a total of 2,816 error instances, covering 18 model responses, and iteratively derived five recurring error types with substantial inter-rater reliability (Fleiss’ $\kappa=0.74$). The remaining cases were then independently labeled using the agreed error taxonomy.
Table  4 summarizes the distribution of error types across model categories. Errors related to over-inference were most prevalent (33.5%), followed by evidence misidentification (23.4%) and hallucination (19.8%). We further observed systematic differences across model types (Figure  5): smaller models exhibited a higher proportion of hallucination errors (48.78%), whereas reasoning-oriented models committed over-inferred more frequently (46.61%) beyond the information explicitly provided in student responses.
Figure 5: Error type distribution by model category.
@@ 6 Discussion
Our findings imply that current LLMs are not yet reliable for cognitive skill diagnosis without intermediate verification, as errors cascade throughout the diagnostic pipeline. Performance was strongly constrained by the evidential strength in student responses rather than by model capabilities. This challenge is closely tied to the nature of students’ handwritten work, which often omits intermediate reasoning and expresses ideas in implicit or non-linear ways, making diagnostic evidence difficult to assess [11, 24]. Under such conditions, LLMs tend to over-attribute evidential strength and, in some cases, arrive at incorrect diagnoses with asserted evidence, raising concerns for real-world educational use.
Our findings point to an important opportunity: richer diagnostic evidence could improve cognitive diagnosis. One promising direction is to design problem-solving tasks and interfaces that better elicit students’ reasoning processes and justifications, help them organize their steps tidily, and prompt reflection on strategy choices. For example, interactive interfaces could track diagnostic evidence in real-time and request clarification when evidence appears weak, shifting from post-hoc inference to active elicitation of cognitive processes.
Beyond task design, our qualitative error analysis highlights the importance of teacher-in-the-loop approaches for mitigating diagnostic failures. Diagnostic errors followed recurring patterns across key stages of the diagnostic process, including identifying evidence, interpreting diagnostic criteria, and synthesizing intermediate reasoning into final judgments. At these stages, teachers can intervene in targeted ways. Errors such as evidence misidentification can be mitigated by having teachers validate whether the identified evidence is actually present in the student’s work. Rubric misinterpretation can be addressed by allowing teachers to check and correct how model interpretations align with diagnostic criteria. For over-inference, teachers can scrutinize whether inferred cognitive skills are sufficiently supported by observable evidence, preventing unwarranted extrapolation. Consistency issues and hallucinations require human oversight to ensure coherence between intermediate reasoning and final diagnostic judgments. Rather than treating model outputs as final assessments, such selective human intervention can reduce diagnostic errors, consistent with prior work on evidence-supported reasoning [28] and human-AI complementary assessment workflows [26]. Overall, responsibly integrating LLMs into educational assessment requires moving beyond fully automated diagnosis toward transparency and selective human verification.
7 Future Work
While we present a novel, carefully crafted benchmark for evaluating LLM-based cognitive skill diagnosis, several directions remain for future work. First, the benchmark can be extended in a more targeted manner. Synthetic data generation [3] could selectively augment diagnostically challenging cases, such as student responses with implicit or vague evidence. Our analysis of 3,036 verdicts highlights where such challenges arise, motivating controlled data expansion rather than indiscriminate scaling. Second, future work can broaden the range of cognitive demands represented in the dataset. Reasoning-oriented problems that require students to generalize or justify their thinking are currently underrepresented in MathCog, and expanding coverage across problem types, topics, and grade levels would enable more comprehensive evaluation. Finally, beyond dataset expansion, future work can explore richer evaluation settings for cognitive diagnosis, such as human–AI collaborative diagnosis scenarios or analyzing how LLMs revise diagnostic judgments when provided with additional evidence. It could shed light on the reliability and practical use of LLM-based diagnostic systems in educational contexts.
Appendix 0.A Appendix
0.A.1 Prompts
0.A.2 System Prompt
# **Task Description**
## **Evaluation Categories**
## **Input Format**
You will receive the following data:
- **Problem**: A math problem given to a student.
- **Answer**: The correct step-by-step solution.
- **Response**: The students response to the problem.
- **Check Items**: A set of specific skills to evaluate.
## **Output Format**
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
0.A.3 User Prompt
# **Task**
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
**Problem:**
There is a two-digit natural number whose tens digit is 11. If the number that changes the tens and ones digits of this natural number is 99 less than 55 times the first number, find the first number.
**Answer:**
If the number in the ones place is $x$, this natural number is $10+x$. The number where the tens digit and the ones digit are swapped is $10x+1$ because the tens digit is $x$ and the ones digit is 11. The changed number is 99 smaller than 55 times the first number, so 10x+1=5(10+x)910x+1=5(10+x)-9 $10x+1=5x+41$ $5x=40$ $x=8$ Therefore, the first number is $18$.
**Response:**
11 $x$ $x$ 11 10x+1=5(10+x)910x+1=5(10+x)-9 10x+1=50+5x910x+1=50+5x-9 $41$ $5x=40$ $x=8$
**Check Items:**
**Output:**
"skills": [
"evidence": "10 x+1=5(10+x)-9",
"verdict": "Evident Yes"
"evidence": "10 x+1=5(10+x)-9",
"verdict": "Evident Yes"
0.A.4 MathCog Dataset Details
This section provides supplementary details of the MathCog dataset, including the diagnostic workflow, cognitive skill taxonomy, inter-rater reliability of teachers’ diagnoses, dataset composition, and representative failure cases of model diagnoses.
@@ 0.A.4 MathCog Dataset Details
Carry out algorithmic procedures for $+,-,\times,\div$ or a combination of these
@@
Topics Content Domain Difficulty
1. Problems involving the digits of numbers Number 1, 3
2. Solving for unknowns under special conditions on the solution Algebra 1, 2, 3
3. Finding unknowns when two equations have the same solution Algebra 1, 2
4. Applying linear inequalities to geometric figures Geometry 2, 3
5. Applying linear inequalities to pricing Algebra 1, 2
7. Using square roots to calculate lengths in geometric figures Geometry 2, 3
8. Performing simple addition and subtraction of square roots Number 1, 2
9. Applying multiplication formulas Number 2, 3
10. Determining coefficients or constants to complete the square Number 1, 2, 3
11. Solving quadratic equations by factoring Algebra 1, 2
12. Rewriting expressions in perfect square form Algebra 1, 2
15. Finding the quadratic function given the vertex and another point on its graph Algebra 1, 2
Table 6: The topics and difficulty levels of problems. This information comes from the original dataset’s metadata, which details the topic and difficulty level (1-3) of each problem. The isomorphic problems require the same mathematical concept to solve, but the difference in numbers makes one more tricky and complicated than the other.
Topics 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
@@ 15. Finding the quadratic function given the vertex and another point on its graph Algebra 1, 2
Table 7: The inter-rater percentage absolute agreement of each topic. The percentage indicates the ratio of unanimous verdicts in each teacher group.
Recall Recognize Classify/Order Compute Determine Represent Implement
Evident Yes 413 434 92 522 441 122 307
Evident No 62 56 8 125 54 33 140
Vague Yes 20 16 2 26 25 0 18
Vague No 35 28 0 20 16 3 18
Student responses 530 534 102 693 536 158 483